Archive for June, 2008

Planned Maintenance: 05th July 2008 03.00hrs – 04.00hrs

Posted: Tuesday, June 24th, 2008 at 12:26 by Neil Watson

During the maintenance window above we will be upgrading the IOS software on wolv-te.core router. The introduction of the new IOS will require a router reload which will impact users connected to our 21CN DSL service at the Wolverhampton node.

It is anticipated that the reload will take no longer than 10 – 15 minutes to complete.

  1. Steve Lalonde says:

    Maintenance completed as designed, traffic is now routed over the optimal path. Please report any issues via the usual support channels.

Planned Maintenance – 28th June 2008 02.00hrs – 06.00hrs

Posted: Tuesday, June 24th, 2008 at 12:22 by Neil Watson

During the maintenance window above we will be upgrading the IOS software on several core routers. The introduction of the new IOS will require a router reload which will impact local services during this window. The routers will be upgraded consecutively to minimise any disruption. The routers affected, and the proposed time of reboot are:

birmingham-te.core – 03.00hrs
Affected customers: This router will impact 21CN ADSL connections on the Birmingham node and customers directly connected to this equipment.

telford-dc1.core – 03.30hrs
Affected customers: All customers connected via DC1 in our Telford HQ will be affected.

bir.core – 04.00hrs
Affected customers: This router will affect all customers who are directly connected and those that connect ‘behind’.

amsterdam.core: 04.30hrs
Affected customers: Peering via Amsterdam will be affected.

frankfurt.core: 05.00hrs
Affected customers: Peering with DECIX via Frankfurt will be affected.

camb.core: 05.30hrs
Affected customers: all customers connected in our Cambridge POP will be affected by the reload.

It is anticipated that the reloads will take no longer than 10 – 15 minutes to complete. We are currently in the process of contacting directly connected customers to advise of this notice.

  1. Steve Lalonde says:

    All maintenance work completed successfully, please report any issues via the usual support channels.

Oubound VoIP Issues

Posted: Friday, June 20th, 2008 at 12:31 by Neil Watson

We are currently experiencing issues with outbound call services via our VoIP platform. We are investigating the cause and will update as soon as possible.

Outage: Telehouse East

Posted: Thursday, June 12th, 2008 at 15:49 by Neil Watson

We are currently experiencing an outage in Telehouse East that is affecting our DSL platform and any permanent connections that terminate in this location. We are working to restore service as soon as possible. During this time we have an increased number of calls which is leading to increased wait times on the phones, we are endeavouring to answer all calls as quickly as possible. Apologies for the problems that this outage will cause.

  1. Neil Watson says:

    Further information: The power for a single rack has tripped and, unusually, the master trip switch has also been triggered. We are currently waiting for a Telehouse engineer to restore power. As soon as we have more information we will update you further.

  2. Neil Watson says:

    Update: We have now been advised that power has been restored. Our engineer is bringing all equipment back in to service sequentially. We will advise as soon as possible that all service has been restored.

  3. Neil Watson says:

    All services are now up and running, however due to a massive imbalance with our central pipes, it is necessary to rebalance them fairly aggressively. We will be doing this as soon as possible to ensure a reasonable balance is achieved for everyone. Once again, apologies.

  4. Neil Watson says:

    A RFO for yesterdays problems is now available.

    Outage Report

LNS Issue

Posted: Monday, June 9th, 2008 at 13:40 by Wayne Hawtin

We have experienced a problem with one of our DSL gateways which caused a momentary drop for some users. The connections affected have re-established but this may result in some imbalance in our centrals. Users experiencing problems should be able to resume normal service by rebooting their router.

Leased Line Packet Loss: redbus-hex.edge.enta.net

Posted: Monday, June 2nd, 2008 at 22:28 by Martyn Haynes

We are currently experiencing severe packet loss with one of the bearers carrying PPC circuits into us, affecting some of our PPC customers. A fault has been raised with our supplier, who intend to investigate – this will involve the removal of this link which will cause some PPC circuits to go down for approx. 20 minutes. Circuits with working ADSL backup should not be affected.

We currently have no ETA for this emergency work but further updates will be posted as they become available.

  1. Martyn Haynes says:

    I have been advised that an engineer is in transit to the site and will proceed testing (which will include the removal of the PPC link) within the hour – further updates will be provided when available.

Change of LNS names

Posted: Monday, June 2nd, 2008 at 18:43 by Neil Watson

The LNS routers are currently known via an customer friendly name e.g. elm and larch, our internal automated systems however use a simplified version. Due to the need for automated updates and reverse DNS, we will be changing the customer facing names to match the internal ones. The new list of names can be found here. Apologies for the short notice.